OpenStudy (imamermaid29):

Help please!! Will fan and medal Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y 0 3 1 5 Line 2 x y –1 1 –2 –1 Line 1: y = ____x + ____ Line 2: y = ____ x + ____

OpenStudy (anonymous):

do you know how to find the slope?

OpenStudy (anonymous):

"no" is a fine answer, just asking

OpenStudy (imamermaid29):

sometimes i have it other times i don't its just really hard

OpenStudy (anonymous):

on the first graph you have two points given \((0,3)\) and \((1,5)\) right

OpenStudy (anonymous):

even though they are given in the form of a table

OpenStudy (anonymous):

how many units right from \(x=0\) to \(x=1\)? (this is not a trick question, the answer is kind of obvious)

OpenStudy (imamermaid29):

1

OpenStudy (anonymous):

right and how many unit up from \(y=3\) to \(y=5\)?

OpenStudy (imamermaid29):

2

OpenStudy (anonymous):

zactly

OpenStudy (anonymous):

right 1, up 2 slope is \(\frac{2}{1}+2\) that is what the slope means how many units up or down in the y direction (vertical) for each single unit right in the x direction (horizontal)

OpenStudy (anonymous):

oops typo there, i meant right 1, up 2 slope is \(\frac{2}{1}=2\)

OpenStudy (anonymous):

right 1, up 2, slope is 2

OpenStudy (imamermaid29):

so the equation would be y= 2x + 2 ??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the first 2 is right, for the slope, but your second 2 is wrong how did you get that?

OpenStudy (imamermaid29):

wait is it 3??

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

the second number, the "\(b\)" in \[y=mx+b\] is the "y intercept" it is the y value when the x value is 0

OpenStudy (anonymous):

when it the corresponding y for \(x=0\)?

OpenStudy (anonymous):

oh sorry i didn't see your second answer yes, it is 3

OpenStudy (anonymous):

if \(x=0\) you have \(y=3\)

OpenStudy (imamermaid29):

:)

OpenStudy (anonymous):

ok now the next one requires more work, but not much more

OpenStudy (imamermaid29):

okay.

OpenStudy (anonymous):

how many units right from \(x=-2\) to \(x=-1\)?

OpenStudy (imamermaid29):

1

OpenStudy (anonymous):

yes, and how many unit up from \(y=-1\) to \(y=1\)?

OpenStudy (imamermaid29):

2

OpenStudy (anonymous):

so guess what the slope is?

OpenStudy (imamermaid29):

1?

OpenStudy (imamermaid29):

wait..

OpenStudy (anonymous):

yeah i am waiting because it is definitely not 1

OpenStudy (imamermaid29):

3

OpenStudy (anonymous):

over one, up 2, slope is ___

OpenStudy (anonymous):

sort of reminiscent of the last one right?

OpenStudy (imamermaid29):

2?

OpenStudy (anonymous):

yes

OpenStudy (imamermaid29):

im just thinking way to hard

OpenStudy (anonymous):

yes for sure, this is not very complicated

OpenStudy (anonymous):

btw we can also get the slope even if the points are not one unit apart using \[\frac{y_2-y_1}{x_2-x_1}\] in this case we would have got \[\frac{1-(-1)}{-1-(-2)}=\frac{2}{1}=2\]

OpenStudy (imamermaid29):

thats how i usually do it

OpenStudy (anonymous):

now for the y intercept

OpenStudy (anonymous):

you want to know what you get when \(x=0\) from -1 to 0 is one unit right, go up another two units in the y from 1 to ?

OpenStudy (imamermaid29):

3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is the same line as the last one

OpenStudy (imamermaid29):

thank you so much!! i am so sorry i wasted your time im pretty slow.

OpenStudy (anonymous):

hey my pleasure many people post and don't even respond, happy to help someone who is learning it is not like you are born knowing this

OpenStudy (imamermaid29):

:) i respond whenever i post a question because i like to work it out so i can actually learn

OpenStudy (imamermaid29):

anyways have a goodnight/day depending where you live :D